Circular No. 5779

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 A. C. Porter, Kitt Peak National Observatory, reports: “Nightly low-

resolution (∼ 0.8 nm) spectroscopic monitoring of SN 1993J in the range 400-750 nm has been carried out on the KPNO 2.1-m telescope (+ Goldcam spectrograph) on Apr. 15-19 UT by J. Stocke, E. Perlman, and E. Ellingson (University of Colorado); on Apr. 20-22 by T. Lauer (KPNO) and M. Postman (Space Telescope Science Institute); and on Apr. 24-30 by F. Owen (National Radio Astronomy Observatory) and M. Ledlow (University of New Mexico and NRAO). An absorption feature in the peak of the broad Hα emission line, which first became visible on ∼ Apr. 19, has grown into a prominent notch ∼ 10 nm wide on Apr. 30. Continued and higher-resolution spectroscopic and polarimetric observations of this feature are urged. Observers are also advised that instrumentation schedules may prevent regular JHK photometry or near-infrared spectroscopy (800-1000 nm, including the calcium triplet) from being done at KPNO during May. Other observatories should make a special effort to fill in these gaps.”

GRS 1915+105 AND HERCULES X-1 S. Brandt, A. J. Castro-Tirado, and N. Lund, Danish Space Research

Institute, report: “The x-ray transient GRS 1915+105 (IAUC 5590, 5619) has been observed by the WATCH experiment on EURECA from Apr. 27.25 to 29.90 UT during the first of four offset pointings of the satellite. After nearly one year, the source is still very bright (580 ± 60 mCrab at 6-12 keV; 160 ± 30 mCrab at 12-20 keV). This strengthens the case for an association between the x-ray transient and the variable radio source recently reported (IAUC 5773). The x-ray source Her X-1 was observed during the same pointing with average fluxes 170 ± 30 mCrab at 6-12 keV and 280 ± 50 mCrab at 12-20 keV.”

HR 1469 = NSV 1671 U. Bastian, Heidelberg, reports that a photograph taken by D. Becker,

Schongau, Germany, shows a star trail of HR 1469 (cf. IAUC 5656, 5675) at its normal brightness from 1992 Nov. 7.906 to 7.917 UT, suggesting that the fading reported on IAUC 5656 (if real) lasted < 3.3 days.

1993 April 30 Daniel W. E. Green